40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 5931     Entry time: Thu Nov 17 21:12:09 2011
Author: Koji 
Type: Update 
Category: PSL 
Subject: HEPA setting changed 

[Koji, Suresh]

8:50PM HEPA@100% for the test

8:55PM HEPA@0%

9:20-35PM HEPA level varies from 0%-50%

9:35PM HEPA@40% and left it running at this level

Nov18 1:40 AM HEPA@80% for a work around the PSL table (by KI)

Nov18 4:35 AM HEPA@40% (by KI)

ELOG V3.1.3-